OpenStudy (anonymous):

Multiply and show your work. (3 - 2i)(4 + i).

sam (.sam.):

Is 'i' imaginary?

sam (.sam.):

Btw just distribute them

sam (.sam.):

(a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

yes the i is imaginary

OpenStudy (whpalmer4):

Then remember that \(i^2 = -1\)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

i need help showing the work i know the answer

OpenStudy (anonymous):

Assuming i is the imaginary unit, use i as a variable instead.

OpenStudy (whpalmer4):

Just multiply like you would ordinarily, for example: \[(1-i)(1+2i) = 1*1 + 1*2i - 1*i -i*2i = 1 + 2i - i -2i^2 = 1 +i -2i^2\]Now remember that \(i^2=-1\) and we get:\[1+i-2i^2=1+i-2(-1)=1+i+2=3+i\]

OpenStudy (anonymous):

Simplify and show you work \[\sqrt{-24}\]